OpenStudy (anonymous):

Need this explained? What are the factored forms of the following expressions? d^2+12d+32 d^2-12d+32 x^2-2-42 d^2-18d+80 x^2+3yx-4y^2

OpenStudy (anonymous):

#1-5

OpenStudy (anonymous):

@TuringTest i only need help with the last one can you help me out?

OpenStudy (turingtest):

x^2+3yx-4y^2 (ax + by)(cx + dy) think of the possibles factors for the coefficient of x and y for x you have either a=c=1 or a=c=-1 for y you have b=1, d=-4 b=-1, d=4 b=2, d=-2 b=-2, d=2 b=4, d=-1 b=-4, d=1 noticing that the cross term has a coefficient of 3 should suggest that b=+/-2, d=+/-2 are unlikely try the other potential factors until they foil out properly

OpenStudy (turingtest):

ruling out the middle two options, as I said, there are only 8 possibilities total. Just exhaust the possibilities.

OpenStudy (anonymous):

oh wow thats a lot to figure out

OpenStudy (turingtest):

not really... (x-4y)(x+y)=? (-x-4y)(-x+y)=? (x+4y)(x-y)=? (-x+4y)(-x-y)=? . . . that's already half the possibilities right there

OpenStudy (anonymous):

so if im doing this right the first number will be x-4?

OpenStudy (turingtest):

no, it's got to have a y in it

OpenStudy (anonymous):

oh sorry for got to put y at the end

OpenStudy (anonymous):

so it would be x-4y?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

then its x+y?

OpenStudy (anonymous):

nope its x-y. thanks for your help man i appreciate it!